# The Cosmic Evolution and Spatial Distribution of Multiphase Gas associated with QSOs

### v1

@@ 4 Summary
Cosmic Evolution: The absorption strength of various elements and ions systematically decreases with increasing redshift, following the trend (see Figure 2 and Figure 3). This behavior indicates a progressive enrichment of metals over cosmic time. At large scale ( Mpc), there is a stronger redshift dependence (), which may be due to the delayed enrichment of metals from small to large scales associated with galaxies.
@@
Ion-Specific Trends: Fe/Mg abundance ratios, gauged by the ratio , decline with increasing redshift (see Figure 6). This trend can be attributed to the vigorous star formation activities at high redshift, along with the delayed Fe enrichment from SNIa in comparison to the Mg production by CCSN. Alternatively, this decline may also indicate a top-heavy IMF in the early universe, which would enhance the production of elements are produced. Additionally, the exhibits a slight increase with , potentially due to the differential dust depletion of Fe.
@@ Acknowledgements
EW thanks the support of the National Science Foundation of China (Nos. 12473008) and the Start-up Fund of the University of Science and Technology of China (No. KY2030000200). HZ acknowledges the supports from National Key R&D Program of China (grant Nos. 2022YFA1602902, 2023YFA1607800, 2023YFA1607804) and the National Natural Science Foundation of China (NSFC; grant Nos. 12120101003 and 12373010). ZCH is supported by NSFC-11903031 and USTC Research Funds of the Double First-Class Initiative YD 3440002001. HYW is supported by the National Natural Science Foundation of China (Nos. 12192224) and CAS Project for Young Scientists in Basic Research, Grant No. YSBR-062. YG receives funding from Scientific Research Fund of Dezhou University (3012304024), Shandong Provincial Natural Science Foundation (ZR2024QA212), and the National Natural Science Foundation of China (NSFC, Nos.12033004 and 12233005). The authors gratefully acknowledge the support of Cyrus Chun Ying Tang Foundations.
@@ Appendix A Comparison with ELGs and LRGs
Appendix B The impact of Energy Level
Figure B presents the equivalent widths for different Fe ii (left panel) and Ni ii (right panel) absorption lines. The horizontal axis displays different absorption lines for each ion, with Fe ii absorption lines on the left and Ni ii lines on the right. The figure illustrates that the values for Fe ii vary significantly with redshift, with trends that are not consistent across all lines. This variability suggests that different absorption lines corresponding to different physical conditions of the gas. Although the equivalent width is theoretically related to the transition probability of the absorption line, the underlying processes are complex and merit further investigation, especially with the anticipated availability of extensive datasets from future DESI observations.

### v2

@@ 4 Summary
Cosmic Evolution: The absorption strength of various elements and ions systematically decreases with increasing redshift, following the trend (Figure 3). This behavior indicates a progressive enrichment of metals over cosmic time. The enrichment rate is closely related to the production mechanisms and ionization states of the elements. Specifically, Fe-peak elements appear to exhibit somewhat stronger evolution than -elements, and ions tracing warm gas tend to show more pronounced evolution than those tracing cold or cool gas. However, the statistical significance of these trends remains limited, and future studies with larger samples are needed to further verify these conclusions.
@@
Ion-Specific Trends: Fe/Mg abundance ratios, gauged by the ratio , decline with increasing redshift but show little sensitivity to distance (see Figure 6). This trend can be attributed to the vigorous star formation activities at high redshift, along with the delayed Fe enrichment from SNIa in comparison to the Mg production by CCSN. Alternatively, this decline may also indicate a top-heavy IMF in the early universe, which would enhance the production of -elements are produced.
@@ Acknowledgements
The authors especially thank the referee for the insightful comments and suggestions, which are crucial to improving this work. EW thanks the support of the National Science Foundation of China (Nos. 12473008) and the Start-up Fund of the University of Science and Technology of China (No. KY2030000200). HZ acknowledges the supports from National Key R&D Program of China (grant Nos. 2022YFA1602902, 2023YFA1607800, 2023YFA1607804) and the National Natural Science Foundation of China (NSFC; grant Nos. 12120101003 and 12373010). ZCH is supported by NSFC-11903031 and USTC Research Funds of the Double First-Class Initiative YD 3440002001. HYW is supported by the National Natural Science Foundation of China (Nos. 12192224) and CAS Project for Young Scientists in Basic Research, Grant No. YSBR-062. YG receives funding from Scientific Research Fund of Dezhou University (3012304024), Shandong Provincial Natural Science Foundation (ZR2024QA212), and the National Natural Science Foundation of China (NSFC, Nos.12033004 and 12233005). The authors gratefully acknowledge the support of Cyrus Chun Ying Tang Foundations.
@@ Appendix A Comparison with ELGs and LRGs
Appendix B The impact of mixed stacking
As our study involves measuring a large number of absorption lines spanning 1206–2900 Å, it is indeed unavoidable that some of these features will fall within the Ly forest region of the background QSO. Therefore, it is important to minimize the confusion in the stacked spectra that could arise from overlap between the regions to the blue and red of Ly. A robust approach, as adopted by Pieri et al. (2014) and Morrison et al. (2024), is to stack the spectra within and outside the Ly forest separately to avoid mixing regions with different physical backgrounds.
The stacked spectra have a high ; thus, rather than increased noise, we are primarily concerned that mixing may introduce a systematic bias in the measurements. In Figure B, we present the PyQSOFit fitting results of background QSO spectra both within and outside the Ly forest, as well as the fraction of spectra at each rest-frame wavelength that are affected by forest contamination. We also test the equivalent width measurements using only spectra outside the forest (). It can be seen that including spectra from the Ly forest may lead to slightly larger equivalent widths, but the difference is minimal. And, this difference could be partly due to differences between the samples. Therefore, such mixed stacking is overall acceptable.
Appendix C The impact of Energy Level
Figure C presents the equivalent widths for different Fe ii (left panel) and Ni ii (right panel) absorption lines. The horizontal axis displays different absorption lines for each ion, with Fe ii absorption lines on the left and Ni ii lines on the right. The figure illustrates that the values for Fe ii vary significantly with redshift, with trends that are not consistent across all lines. This variability suggests that different absorption lines corresponding to different physical conditions of the gas. Although the equivalent width is theoretically related to the transition probability of the absorption line, the underlying processes are complex and merit further investigation, especially with the anticipated availability of extensive datasets from future DESI observations.